OpenStudy (anonymous):

What is the third term in the binomial expansion (x – y^5 )^3?

OpenStudy (anonymous):

(a+b)^3 = a^3 + 3a^2b + 3ab^2 + b^3. In the given equation, a = x and b = -y^5.

OpenStudy (anonymous):

Find 3rd term = 3ab^2